# HetGCoT: Heterogeneous Graph-Enhanced Chain-of-Thought LLM Reasoning for Academic Question Answering

### v2

@@ 1 Introduction
Academic question answering in heterogeneous scholarly networks presents essential challenges in integrating structural knowledge with semantic understanding. QA tasks regarding publishing venue selection, paper authorship, and scientific collaboration all require systems to reason over complex networks of papers, authors, venues, and organizations while providing interpretable explanations Shi et al. (2019); Wang et al. (2022).
@@
Current approaches to academic question answering have attempted to address these challenges through various strategies. However, existing methods face significant limitations in addressing these challenges holistically. Heterogeneous graph neural networks (HGNNs) can effectively model complex academic networks Hu et al. (2020a), but struggle with: (1) adapting their representations to different query types and relationship patterns, (2) generating task-specific subgraph selections, and (3) producing natural language explanations for diverse academic QA scenarios. LLMs demonstrate strong semantic understanding Chowdhery et al. (2022) but cannot directly process the rich structural information embedded in academic networks. Existing integration attempts typically focus on single tasks or treat graph information as auxiliary features through simple concatenation, failing to systematically incorporate structural patterns into the reasoning process across diverse academic QA scenarios Zhao et al. (2023).
To address these limitations, we propose HetGCoT (Heterogeneous Graph-Enhanced Chain-of-Thought), a framework that integrates heterogeneous graph neural networks with large language models for academic question answering. HetGCoT transforms graph structural patterns into confidence-weighted natural language reasoning chains through metapath naturalization, enabling LLMs to process complex academic relationships. The framework employs adaptive metapath selection using Heterogeneous Graph Transformer (HGT) Hu et al. (2020b) embeddings and FastGTN-learned Yao et al. (2021) importance weights to dynamically identify task-relevant subgraphs. Through a multistep chain-of-thought reasoning process, HetGCoT anchors on three task-driven analytical foci: analyzing venue patterns for journal recommendation, temporal relationships for authorship queries, and collaboration networks for collaboration discovery. This integrated approach enables deep reasoning-level fusion of graph structures with language understanding across diverse academic QA scenarios.
@@ 2.1 LLMs and Reasoning
LLMs have revolutionized natural language processing through their sophisticated understanding and generation capabilities. Building upon the Transformer architecture (Vaswani et al., 2017), prominent models including GPT (Brown et al., 2020), LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and PaLM (Chowdhery et al., 2022) have achieved remarkable performance across diverse language tasks.
@@ 2.2 Integration of GNNs and LLMs
Retrieval-Augmented Approaches: PathRAG (Chen et al., 2025) enhances LLMs through graph-based retrieval using relational paths, while GNN-RAG (Mavromatis and Karypis, 2024) combines graph neural retrieval with language model reasoning. Generate-on-Graph (Chen et al., 2024) treats LLMs as both agents and knowledge graphs for incomplete QA tasks.
Heterogeneous Graph and Metapath Methods: For academic networks specifically, heterogeneous graph neural networks like HGT (Hu et al., 2020a) and HAN (Wang et al., 2019) model complex relationships between different entity types. Metapath-based techniques provide interpretable relationship modeling through typed connection sequences. Recent work such as Metapath of Thoughts (Solanki et al., 2024) verbalizes metapaths as contextual augmentation for LLMs. While these methods show promise, they typically focus on single tasks or treat graph information as auxiliary features rather than achieving deep reasoning-level integration.
@@ 4.1 Experimental Setup
We evaluate the proposed HetGCoT framework on two academic datasets, OpenAlex and DBLP. To ensure paper quality and the interpretability of results, we extracted a random subgraph from OpenAlex limited to journals ranked "A" or higher according to the CORE list (a widely used venue ranking system for computing research primarily in Australia and New Zealand), while for DBLP, we randomly sampled a subgraph from the entire dataset. We train the models separately on these two datasets. Table 1 summarises the statistics of the heterogeneous graphs extracted for our experiments. For each node type, we retain the following attributes: venue (type, name); paper (type, keywords, abstract, citations, FWCI (field-weighted citation impact), title, year); author (type, organization, name). All textual fields are encoded using Sentence-BERT, yielding initial node representations that capture semantic content in titles, abstracts, and keywords.
@@
We report four metrics: Hit (%)—measures whether any of the true answers are found in the generated response, which is typically employed when evaluating LLMs; H@1 (%)—the accuracy of the top/first predicted answer; F1 (%)—harmonic mean of precision and recall; and NDCG (%)—which weights higher-ranked correct answers more heavily.
@@ 4.2 Baseline Methods
    GAT (Veličković et al., 2018): Graph attention network that captures relative importance between nodes through attention mechanisms
    HGT (Hu et al., 2020a): Heterogeneous Graph Transformer, an architecture designed specifically for heterogeneous graphs
@@
We further conduct ablation studies with different LLM sizes: Qwen-2.5 1.5B, Qwen-2.5 7B, LLaMA-2 7B, LLaMA-2 13B, and LLaMA-3 8B, assessing the framework’s adaptability to varying foundation-model capacities.
@@ 4.3 Results
Table 4: Performance comparison of different-sized LLMs within the HetGCoT framework.
@@ 4.4 Ablation Study
The ablation study reveals the importance of each component in our framework. Comparing the zero-shot baseline with the HGT+CoT variant shows that incorporating chain-of-thought reasoning yields substantial performance gains. Further analysis of individual reasoning steps indicates that each contributes to the final performance, with step 2 (collaboration relationships) showing the most impact when removed. The complete HetGCoT framework outperforms all partial configurations, indicating that the four-step reasoning process works synergistically.
@@ A.1 Experiment Setup Detail
For HGT, We feed the 770-dimensional feature into a two-layer Heterogeneous Graph Transformer (HGT) with type-specific Q-K-V projections, 8 attention heads, and a hidden size of $d=387$. The model is trained for 100 epochs using Adam (learning rate $1\times 10^{-3}$) on the paper–venue link-prediction task. Final per-node embeddings ($h_{i}\in\mathbb{R}^{387}$) are saved for downstream use.
For TastGTN, a lightweight FastGTN autoencoder (7 layers, 8 channels, hidden size 64) is trained to reconstruct the frozen HGT embeddings (minimizing MSE$(Z_{\text{paper}},h_{\text{paper}})$) over 50 epochs with Adam (learning rate $3\times 10^{-3}$) on GPU. After training, we extract the learned relation-importance weights and score each candidate metapath by summing its edge weights.
A.2 Ablation Study
@@ A.3 Algorithm
2:  $H\leftarrow\text{HGT-ENCODE}(G)$triangleright Heterogeneous graph construction
@@
5:  $\text{MP}_{\text{candidates}}\leftarrow\text{GENERATE-METAPATHS}(G,S_{q})$triangleright HGT Metapath Pool Generation
@@
7:   $\text{scores}[\text{mp}]\leftarrow\sum_{(u,v)\in\text{mp}}W[\psi(u,v)]/|\text{mp}|^{\gamma}$triangleright Adaptive Metapaths Selection
@@
[…]
@@ A.4 Prompt Templates
You are an academic graph-reasoning assistant. Your task is to analyze academic collaboration networks to predict which researchers are most likely to be collaborators of a specific author. Please strictly follow the provided reasoning steps and use the provided graph structure information and paper context to make your predictions.

### v3

@@ 1 Introduction
Academic question answering in heterogeneous scholarly networks presents essential challenges in integrating structural knowledge with semantic understanding. QA tasks regarding publishing venue selection, paper authorship, and scientific collaboration all require systems to reason over complex networks of papers, authors, venues, and organizations while providing interpretable explanations Shi et al. (2019); Ji et al. (2021).
@@
Current approaches to academic question answering have attempted to address these challenges through various strategies. However, existing methods face significant limitations in addressing these challenges holistically. Heterogeneous graph neural networks (HGNNs) can effectively model complex academic networks Hu et al. (2020), but struggle with: (1) adapting their representations to different query types and relationship patterns, (2) generating task-specific subgraph selections, and (3) producing natural language explanations for diverse academic QA scenarios. LLMs demonstrate strong semantic understanding Chowdhery et al. (2023) but cannot directly process the rich structural information embedded in academic networks. Existing integration attempts typically focus on single tasks or treat graph information as auxiliary features through simple concatenation, failing to systematically incorporate structural patterns into the reasoning process across diverse academic QA scenarios Pan et al. (2024).
To address these limitations, we propose HetGCoT (Heterogeneous Graph-Enhanced Chain-of-Thought), a framework that integrates heterogeneous graph neural networks with large language models for academic question answering. HetGCoT transforms graph structural patterns into confidence-weighted natural language reasoning chains through metapath naturalization, enabling LLMs to process complex academic relationships. The framework employs adaptive metapath selection using Heterogeneous Graph Transformer (HGT) Hu et al. (2020) embeddings and FastGTN-learned Yun et al. (2022) importance weights to dynamically identify task-relevant subgraphs. Through a multistep chain-of-thought reasoning process, HetGCoT anchors on three task-driven analytical foci: analyzing venue patterns for journal recommendation, temporal relationships for authorship queries, and collaboration networks for collaboration discovery. This integrated approach enables deep reasoning-level fusion of graph structures with language understanding across diverse academic QA scenarios.
@@ 2.1 LLMs and Reasoning
LLMs have revolutionized natural language processing through their sophisticated understanding and generation capabilities. Building upon the Transformer architecture (Vaswani et al., 2017), prominent models including GPT (Brown et al., 2020), LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and PaLM (Chowdhery et al., 2023) have achieved remarkable performance across diverse language tasks.
@@ 2.2 Integration of GNNs and LLMs
Retrieval-Augmented Approaches: PathRAG (Chen et al., 2025) enhances LLMs through graph-based retrieval using relational paths, while GNN-RAG (Mavromatis and Karypis, 2024) combines graph neural retrieval with language model reasoning. Generate-on-Graph (Xu et al., 2024) treats LLMs as both agents and knowledge graphs for incomplete QA tasks.
Heterogeneous Graph and Metapath Methods: For academic networks specifically, heterogeneous graph neural networks like HGT (Hu et al., 2020) and HAN (Wang et al., 2019) model complex relationships between different entity types. Metapath-based techniques provide interpretable relationship modeling through typed connection sequences. Recent work such as Metapath of Thoughts (Solanki et al., 2024) verbalizes metapaths as contextual augmentation for LLMs. While these methods show promise, they typically focus on single tasks or treat graph information as auxiliary features rather than achieving deep reasoning-level integration.
@@ 4.1 Experimental Setup
We evaluate the proposed HetGCoT framework on two academic datasets, OpenAlex and DBLP. To ensure paper quality and the interpretability of results, we extracted a random subgraph from OpenAlex limited to journals ranked "A" or higher according to the CORE list (a widely used venue ranking system for computing research primarily in Australia and New Zealand), while for DBLP, we randomly sampled a subgraph from the entire dataset. We train the models separately on these two datasets. Table 1 summarises the statistics of the heterogeneous graphs extracted for our experiments. For each node type, we retain the following attributes: venue (type, name); paper (type, keywords, abstract, citations, FWCI (field-weighted citation impact), title, year); author (type, organization, name). All textual fields are encoded using Sentence‑BERT, yielding initial node representations that capture semantic content in titles, abstracts, and keywords.
@@
We report four metrics: Hit (%)—measures whether any of the true answers are found in the generated response, which is typically employed when evaluating LLMs; H@1 (%)—the accuracy of the top/first predicted answer; F1 (%)—harmonic mean of precision and recall; and NDCG (%)—which weights higher‑ranked correct answers more heavily.
@@ 4.2 Baseline Methods
    GAT (Veličković et al., 2017): Graph attention network that captures relative importance between nodes through attention mechanisms
    HGT (Hu et al., 2020): Heterogeneous Graph Transformer, an architecture designed specifically for heterogeneous graphs
@@
    HiGPT (Tang et al., 2024): Heterogeneous graph language model for graph-based reasoning
@@
We further conduct ablation studies with different LLM sizes: Qwen-2.5 1.5B, Qwen-2.5 7B, LLaMA‑2 7B, LLaMA‑2 13B, and LLaMA‑3 8B, assessing the framework’s adaptability to varying foundation‑model capacities.
@@ 4.3 Results
Table 4: Performance comparison of different‑sized LLMs within the HetGCoT framework.
@@ 4.4 Ablation Study
The ablation study reveals the importance of each component in our framework. Comparing the zero-shot baseline with the HGT+CoT variant shows that incorporating chain-of-thought reasoning yields substantial performance gains. Further analysis of individual reasoning steps indicates that each contributes to the final performance, with step 3 (collaboration relationships) showing the most impact when removed. The complete HetGCoT framework outperforms all partial configurations, indicating that the four-step reasoning process works synergistically.
@@ 5 Conclusion
Acknowledgments
This work was supported by the Commonwealth Scientific and Industrial Research Organization (CSIRO), Australia, in conjunction with the National Science Foundation (NSF) of the United States, under grant CSIRO-NSF #2303037.
@@ A.1 Experiment Setup Detail
For HGT, We feed the 770-dimensional feature into a two-layer Heterogeneous Graph Transformer (HGT) with type-specific Q-K-V projections, 8 attention heads, and a hidden size of $d=387$. The model is trained for 100 epochs using Adam (learning rate $1\times 10^{-3}$) on the paper–venue link‐prediction task. Final per‐node embeddings ($h_{i}\in\mathbb{R}^{387}$) are saved for downstream use.
[…]
A.2 Statistical Robustness Analysis
To verify the statistical robustness of our approach, we conducted three independent runs with different random seeds on both datasets. Table 6 presents the detailed results across all metrics.
Table 6: Statistical robustness analysis across three independent runs with different random seeds.
Dataset Run Hit H@1 F1 NDCG
OpenAlex Run 1 95.73 91.45 82.91 91.29
Run 2 96.48 92.21 79.90 91.21
Run 3 96.50 92.00 78.00 92.00
[…]
DBLP Run 1 85.31 83.70 64.55 83.49
Run 2 84.70 83.60 64.66 83.39
Run 3 85.27 83.75 64.38 83.46
[…]
The results demonstrate consistent performance across multiple runs. This consistency confirms the stability of our approach across different experimental conditions.
A.3 Evaluation with Reasoning-Type LLMs
To assess the compatibility of our framework with advanced reasoning-type language models, we conducted additional experiments using OpenAI o3 and DeepSeek-R1 without fine-tuning. Table 7 presents the comparative results.
Table 7: Performance comparison with reasoning-type LLMs.
Dataset Model Hit H@1 F1 NDCG
OpenAlex o3 74.82 59.46 39.56 65.38
o3+HetGCoT 90.90 86.18 68.75 85.46
DeepSeek-R1 73.58 60.38 41.51 62.26
DeepSeek-R1+HetGCoT 92.16 88.24 74.51 78.43
HetGCoT (ours) 96.48 92.21 79.90 91.29
DBLP o3 62.58 51.67 41.60 60.08
o3+HetGCoT 78.14 68.62 53.54 67.83
DeepSeek-R1 63.16 57.89 43.86 59.65
DeepSeek-R1+HetGCoT 79.60 71.43 61.22 69.39
HetGCoT (ours) 85.31 83.70 64.55 83.49
The results demonstrate that our HetGCoT framework maintains its effectiveness when applied to reasoning-type language models, with consistent performance improvements observed across different model architectures.
A.4 Ablation Study
@@ A.5 Algorithm
2:  $H\leftarrow\text{HGT-ENCODE}(G)$ $\triangleright$ Heterogeneous graph construction
@@
5:  $\text{MP}_{\text{candidates}}\leftarrow\text{GENERATE-METAPATHS}(G,S_{q})$ $\triangleright$ HGT Metapath Pool Generation
@@
7:   $\text{scores}[\text{mp}]\leftarrow\sum_{(u,v)\in\text{mp}}W[\psi(u,v)]/|\text{mp}|^{\gamma}$ $\triangleright$ Adaptive Metapaths Selection
@@
13:  $\text{prompt}\leftarrow\text{CONSTRUCT-PROMPT}(q,\text{contexts})$
14:  $\text{answer}\leftarrow\text{LLM-COT-REASONING}(\text{prompt})$ $\triangleright$ 4-step reasoning
@@ A.6 Prompt Templates
You are an academic graph‐reasoning assistant. Your task is to analyze academic collaboration networks to predict which researchers are most likely to be collaborators of a specific author. Please strictly follow the provided reasoning steps and use the provided graph structure information and paper context to make your predictions.